Given that the equation of tangent at the point (at^2 , 2at) on the parabola y^2 = 4ax is x - ty + at^2 = 0
A(at1^2 , 2at1) and B(at2^2 , 2at2) are points on this parabola. The equation of the chord AB is 2x - (t1 + t2)y + 2at1at2 = 0
The tangents A and B meet at the point P. Find the coordinates of P.
The line through P parallel to the axis of the parabola meets the chord AB at M. Find the coordinates of M. Prove that M is the midpoint of AB.
1
Expert's answer
2021-02-02T01:44:47-0500

"x=\frac{y^2}{4a},"

"k=x_y'=\frac{y}{2a},"

"x-at_i^2=\frac{2at_i}{2a}(y-2at_i),"

"\begin{cases}
 x-at_1^2=t_1(y-2at_1), \\
 x-at_2^2=t_2(y-2at_2);
\end{cases}"

"\begin{cases}
x_0=at_1t_2, \\
 y_0=a(t_1+t_2); 
\end{cases}" – coordinates of point P.

Line PM (parallel to the x-axis) has an equation "y=a(t_1+t_2)."

Midpoint of AB has coordinates "(\frac a2(t_1^2+t_2^2),~a(t_1+t_2))."

AB equation: "2x-(t_1+t_2)y+2at_1t_2=0,"

with a PM equation we'll have

"\begin{cases}
x_0=\frac a2(t_1^2+t_2^2), \\
 y_0=a(t_1+t_2); 
\end{cases}" – coordinates of point M, and they are equal to the coordinates of point P.